Document Actions

 

Status of Systems

Note: Systems may not be available during the reserved maintenance and testing time frame, Wednesdays from 6:00 a.m. to 7:00 a.m.

Current Status of Systems

No outages reported.


Ongoing

Beware of Emails Asking For Your Information!
Many people are receiving 'phishing' emails which ask you for sensitive/personal info: (e.g. login ID's and passwords, date of birth, Social Security, etc.). Simon's Rock would NOT ask you for this or other personal information. *NEVER REPLY* or click on links in such emails. Doing so can compromise your personal information, cause your email account and login ID to be suspended, and even result in identity theft. Please contact ITS at 413-528-7371 with concerns or for additional information.


Outages

Januray 11, 2012

6:00 AM - Email server has been successfully migrated. Everything appears to be working as normal

3:00 AM - Mail server has been taken offline to migrate all email services to a new virtual server. Will update again when email services are restored.

May 17, 2011

Unexpected hard disk failure on an important server caused email and any authenticated services to fail between about 4:55PM and 6:20PM

May 9, 2011

Our primary internet connection went down, causing us to fail over to our backup connection. This did affect Banner connectivity along with other services

March 21, 2011

Very early in the morning (2:18 AM), an on-campus misconfiguration began to manifest itself. No problems were reported until after 6 AM. By 7AM, the lack of access to resources from on-campus locations was confirmed. Troubleshooting commenced as ITS staff arrived. Off-campus connectivity was only briefly interrupted (as attempts were made to remedy the on-campus situation). Around 10AM a problem was identified and corrected, and all services were restored shortly thereafter.

Feb. 19, 2011

Power outage commenced at some time after 1PM. Servers remained up for several hours on UPS backup power, but that gave out before the power was restored (around 6:00PM)

Feb. 16, 2010

Scheduled network maintenance at Fisher. Intermittent outages from 11 AM to 1 PM.

Feb. 12, 2010

Student Union network connection was down from sometime overnight until 9:20 AM, due to UPS problem at Student Union switch rack.

Feb. 10-11, 2010

Students were unable to login to the Inside website from approximately 3 p.m. on Feb. 10 until noon on Feb. 11, due to an LDAP issue.

Feb 5, 2010

Network outage to Orchard 1 + 2, due to failed media converter.

Feb 1, 2010

Network outage for faculty and printers at Fisher, due to switch configuration error.

October 8, 2009

Possible Massive Upstream / Service Provider Outage.
552975: Oct  8 16:31:27.059 EDT: %LINK-3-UPDOWN: Interface Serial0/1/0, changed state to down
552976: Oct  8 16:31:27.223 EDT: %LINK-3-UPDOWN: Interface Serial0/0/0, changed state to down
552977: Oct  8 16:31:27.227 EDT: %LINK-3-UPDOWN: Interface Multilink1, changed state to down
552978: Oct  8 16:31:28.059 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial0/1/0, changed state to down
552979: Oct  8 16:31:28.223 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial0/0/0, changed state to down
552980: Oct  8 16:31:28.227 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Multilink1, changed state to down
552985: Oct  8 16:32:24.717 EDT: %CONTROLLER-5-UPDOWN: Controller T3 1/0, changed state to down
552986: Oct  8 16:32:26.717 EDT: %LINK-3-UPDOWN: Interface Serial1/0, changed state to down
552987: Oct  8 16:32:27.717 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial1/0, changed state to down

Sept 30, 2009

At sometime around 9:40AM local time ITS began receiving calls about partial outages.   Troubleshooting commenced and by 10:17, we'd restored service.   Affected services came back up without a problem once we'd correctly diagnosed and remedied the problem.  We expect no aftereffects (other than some email trickling in a bit more slowly than it might have arrived without any problems).

April 29, 2009

29 1122: All Internet connectivity issues pertaining to the previous outage should now be resolved (including student networks to Internet).

WED APR 29 0043: DS3 RESTORED

WED APR 29 0030: T3 DOWN:
078000: Apr 29 00:10:24.324 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial1/0, changed state to down
078001: Apr 29 00:12:05.699 EDT: %BGP-5-ADJCHANGE: neighbor 64.213.79.249 Down BGP Notification sent
078002: Apr 29 00:12:05.699 EDT: %BGP-3-NOTIFICATION: sent to neighbor 64.213.79.249 4/0 (hold time expired) 0 bytes
078010: .Apr 29 00:29:07.539 EDT: %LINK-5-CHANGED: Interface Serial1/0, changed state to administratively down
078011: .Apr 29 00:29:11.227 EDT: %LINK-3-UPDOWN: Interface Serial1/0, changed state to up
078012: .Apr 29 00:29:12.227 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial1/0, changed state to up
078014: .Apr 29 00:29:34.355 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial1/0, changed state to down
078016: .Apr 29 00:43:24.378 EDT: %LINEPROTO-5-UPDOWN: Line protocol on Interface Serial1/0, changed state to up
078017: Apr 29 00:43:42.623 EDT: %BGP-5-ADJCHANGE: neighbor 64.213.79.249 Up

April 3, 2009 10:15 AM

Reports of network outages in the Student Union, Blodgett, and the Daniel Arts Center. ITS is just beginning investigation. Cause is not yet known, thus no estimate of time for restoration of service.

January 15, 2009, 10:28 a.m.

Webmail was down for about two hours. (rm)

Nov 30, 2008

Network outage affecting
Foster House/Mods, Physical Plant, Orchard Houses, Gatehouse caused by fiber cut. Resolved 1600 Dec. 1. (pcl)

Nov 26, 2008 01:00 AM

Wireless authentication server migration completed. Other upgrades completed and servers rebooted. All systems should be online. (pcl)

Nov 14, 2008 7:35-7:44 AM

This was an "outside of the regular maintenance window" upgrade event that we announced on the previous afternoon. After performing this quick upgrade, we should now be able to tolerate a longer power outage (and still keep this main campus website running).

Services Affected: Authentication/Directory services, DNS services (cderr)

Aug 28, 2008 8:45 AM - 10:00 AM


We lost campus power sometime after 8:30AM and by 10:00 the generator was up and running (and we had brought all systems back on-line shortly thereafter). We also shut down non-essential services at 1:00PM (leaving email and web access available) in preparation for switching back over to power from outside (so that the generator could be turned off). (cderr)

Aug 11, 2008 3:30 AM - 12:30 PM

Hardware failure in a critical server meant that our LDAP system became unavailble, breaking many services on campus (the most critical being email). A new server was quickly reprovisioned and backups were used to instantiate a new LDAP server. (cderr)

May 9, 2008 10:18-10:27 AM

Our Cisco router hiccuped and took down our off-campus link for a few minutes. (cderr)

May 4, 2008 around 8:00 p.m.

After working with Cisco tech support to fix an issue with our router, the campus is back online. The length of this outage means that over the 24 hours following, email may still come trickling in. (cderr)

May 4, 2008 -- around 4:00 p.m .our upstream link was restored for good.

We are notified that everything is good with our upstream link and begin troubleshooting our end. (cderr)

May 3, 2008 -- sometime before 11:45 a.m. the upstream connection to the internet went down.

The situation was a major one impacting a major optical fiber link from Pittsfield to Springfield. This was a network outage "upstream" that affected many various other bandwidth consumers (not just Simon's Rock). During the outage, resources on campus remained operational (on campus email still worked for example but not email coming onto or leaving campus). (cderr)

February 13th, 2008 8:45 a.m. - 11:45 a.m.

We lose AC power for several hours, taking down our network connection. (cderr)

February 6th, 2008 -- 12:36 p.m.

Back online. (cderr)

February 6th, 2008 -- 11:34 a.m.

Same Boston-Springfield link goes down again upstream. (cderr)

February 5th, 2008 -- 7:04 p.m.

Link is restored and we again have upstream connectivity. (cderr)

February 5th, 2008 -- 6:01 p.m.

An upstream fiber link between Springfield and Boston goes down. (cderr)

System Changes

Feb 24, 2009 Please read retiring minerva to learn about the pending retirement of one of our longest-running (and most used/loved) webservers.

Feb 23, 2009.  Upgraded pepper to current platform.  This required a reboot.  As far as we know, everything is functioning properly.  Please report problems to sysalert@simons-rock.edu if you find anything amiss.

February 5, 2009.  The weekly updates which were applied overnight introduced configuration changes which broke wireless authentication.   Fixed by 9:05AM

January 6, 2009. The network optical fiber connections that go through the library were upgraded the week of Tuesday, January 6th. Consequently network connections in the library, Crosby, the classroom buildings, the lecture center and ITS offices were down for all or part of those days.

January 7, 2009. ITS will be shutting down the primary mail storage server on Wednesday Jan. 7 at 10 PM in order to migrate user mail boxes to a new server. (pcl)

This migration is required as part of a systems upgrade to increase mail delivery reliability, offer increased uptime during short power outages, improve mailbox recovery viability, and enable better mailbox storage management. The work is expected to be completed no later than 7 AM the next day, Thursday, Jan. 8.

What this means for you, is that starting at approximately 10 PM on Wednesday Jan. 7, you will be unable to access your Simon's Rock e-mail, including access via Webmail, Thunderbird, Mac Mail, and/or Outlook. Mailing lists hosted at Simon's Rock will be nonfunctional.

However, no mail will be lost during this period, and any incoming mail from the Internet will just be queued on our inbound perimeter server. We hope to complete this migration well before 7 AM the next day.

Thank you for your cooperation. If you have any questions, feel free to email us at its@simons-rock.edu.


UPDATE Thurs. January 8, 2009 The mailserver has been upgraded and all mail services appear to be up. Users should not expect to see any changes in the way they access or send mail. If you do have trouble accessing your Simon's Rock email today, your primary point of contact at ITS will be Charlie Derr at (413) 528-7344. Thank you. (PCL)